# Swiss Parliaments Corpus Re-Imagined (SPC_R): Enhanced Transcription with RAG-based Correction and Predicted BLEU

### v1

@@ Abstract
This paper presents a new long-form release of the Swiss Parliaments Corpus, converting entire multi-hour Swiss German debate sessions (each aligned with the official session protocols) into high-quality speech–text pairs. Our pipeline starts by transcribing all session audio into Standard German using Whisper Large-v3 under high-compute settings. We then apply a two-step GPT-4o correction process: first, GPT-4o ingests the raw Whisper output alongside the official protocols to refine misrecognitions, mainly named entities. Second, a separate GPT-4o pass evaluates each refined segment for semantic completeness. We filter out any segments whose Predicted BLEU score (derived from Whisper’s average token log-probability) and GPT-4o evaluation score fall below a certain threshold. The final corpus contains 801 hours of audio, of which 751 hours pass our quality control. Compared to the original sentence level SPC release, our long-form dataset achieves a 6-point BLEU improvement, demonstrating the power of combining robust ASR, LLM-based correction, and data-driven filtering for low-resource, domain-specific speech corpora.
@@ 1 Introduction
  • A two-step large language model (LLM) approach in which a first model corrects the transcription and a second, independent model evaluates that correction.
@@ 3 Transcription with Whisper Large-v3
Our transcription pipeline uses Whisper Large-v3 via WhisperX Bain et al. (2023) under high-compute settings, namely beam_size set to 10, best_of set to 10, and log_prob_threshold set to –2. All transcriptions are performed on an NVIDIA A4500 GPU with 20 GB of VRAM, using float16 precision and a batch_size of 8. These high-compute settings further improve results, as shown in Figure 5. For all transcribed parliament sessions, we store Whisper’s avg_log_prob output, which reflects the model’s prediction confidence and exhibits strong predictive power for transcription quality, as described in Subsection 3.1.
@@ 3.1 BLEU Prediction
Figure 3: Distribution of Predicted BLEU scores across SPC_R ($N$ = 131’291 data segments).
@@
Hence, the Predicted BLEU score derived from Whisper’s avg_log_prob can be used to identify and select high-quality transcription segments (see Section 5).
@@ 4 Transcript correction using GPT-4o
    Correction Stage: GPT-4o is used to refine the initial transcription by prompting it to correct errors, segment by segment. Corrections are based on information injected from the official manual summaries of the parliament session corresponding to the audio segment using Retrieval-Augmented Generation (RAG, see Subsection 4.1).
@@ 4.2 Correction Stage
Figure 5: Word Error Rates (WER) for Whisper Large-v3 under three configurations: standard settings, after applying GPT-4o correction, and using high-compute settings (enhanced settings) with GPT-4o correction.
@@ 5 Selecting Data and Train/Test Split
Figure 7: Logic used to build high-quality SPC_R corpus dataset. Size of initial dataset "Data" is 801 hours of audio, size of high-quality dataset "SPC_R" is 751 hours.
@@
This leads to a high quality corpus of 751 hours of Swiss German audio with paired Standard German transcriptions. For the test set, 50 hours are selected with at least a BLEU score of 70 and segments being evaluated as category 3 (as described in Section 4.3). The train/test split is therefore 701/50 hours.
@@ 7 Conclusion
We present SPC_R, transcribed with Whisper Large-v3 on high-compute settings, corrected with context by GPT-4o, and evaluated for quality by GPT-4o-mini. This process results in a corpus of 751 hours of high-quality spoken Swiss German paired with Standard German text.

### v2

@@ Abstract
This paper presents a new long‑form release of the Swiss Parliaments Corpus, converting entire multi‑hour Swiss German debate sessions (each aligned with the official session protocols) into high‑quality speech–text pairs. Our pipeline starts by transcribing all session audio into Standard German using Whisper Large‑v3 under high‑compute settings. We then apply a two‑step GPT‑4o correction process: first, GPT‑4o ingests the raw Whisper output alongside the official protocols to refine misrecognitions, mainly named entities. Second, a separate GPT‑4o pass evaluates each refined segment for semantic completeness. We filter out any segments whose Predicted BLEU score (derived from Whisper’s average token log‑probability) and GPT‑4o evaluation score fall below a certain threshold. The final corpus contains 801 hours of audio, of which 555 hours pass our quality control. Compared to the original sentence level SPC release, our long‑form dataset achieves a 6‑point BLEU improvement, demonstrating the power of combining robust ASR, LLM‑based correction, and data‑driven filtering for low‑resource, domain‑specific speech corpora.
@@ 1 Introduction
  • A two‑step large language model (LLM) approach in which a first model corrects the transcription and a second, independent model evaluates that correction.
@@ 3 Transcription with Whisper Large-v3
Our transcription pipeline uses Whisper Large‑v3 via WhisperX Bain et al. (2023) under high‑compute settings, namely beam_size set to 10, best_of set to 10, and log_prob_threshold set to –2. All transcriptions are performed on an NVIDIA A4500 GPU with 20 GB of VRAM, using float16 precision and a batch_size of 8. These high‑compute settings further improve results, as shown in Figure 5. For all transcribed parliament sessions, we store Whisper’s avg_log_prob output, which reflects the model’s prediction confidence and exhibits strong predictive power for transcription quality, as described in Subsection 3.1.
@@ 3.1 BLEU Prediction
Figure 3: Distribution of Predicted BLEU scores across SPC_R (NN = 131’291 data segments).
@@
Hence, the Predicted BLEU score derived from Whisper’s avg_log_prob can be used to identify and select high‑quality transcription segments (see Section 5).
@@ 4 Transcript correction using GPT-4o
    Correction Stage: GPT-4o is used to refine the initial transcription by prompting it to correct errors, segment by segment. Corrections are based on information injected from the official manual summaries of the parliament session corresponding to the audio segment using Retrieval‑Augmented Generation (RAG, see Subsection 4.1).
@@ 4.2 Correction Stage
Figure 5: Word Error Rates (WER) for Whisper Large‑v3 under three configurations: standard settings, after applying GPT‑4o correction, and using high-compute settings (enhanced settings) with GPT‑4o correction.
@@ 5 Selecting Data and Train/Test Split
Figure 7: Logic used to build high-quality SPC_R corpus dataset. Size of initial dataset "Data" is 801 hours of audio, size of high-quality dataset "SPC_R" is 555 hours.
@@
This leads to a high quality corpus of 555 hours of Swiss German audio with paired Standard German transcriptions. For the test set, 30 hours are selected with at least a BLEU score of 70 and segments being evaluated as category 3 (as described in Section 4.3). The train/test split is therefore 525/30 hours.
@@ 7 Conclusion
We present SPC_R, transcribed with Whisper Large-v3 on high-compute settings, corrected with context by GPT-4o, and evaluated for quality by GPT-4o-mini. This process results in a corpus of 555 hours of high-quality spoken Swiss German paired with Standard German text.